# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ A.1 Proofs of results in Section 4
Applying the expansion $(1-x)^{-1}=1+x+x^{2}+O(x^{3})$ with $x=\frac{4}{d}+\epsilon_{m}$ shows (41). Similarly, Theorem A.1, applying the expansion $(1-x)^{-1}=1+x+x^{2}+O(x^{3})$ in the denominator of $\frac{1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}}{1-\frac{2}{\sqrt{d}}-\epsilon_{m}^{\prime}}$, and multiplying by the numerator $1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}$ give (42). ∎
A.3 Proofs of Proposition 3.1, Proposition 3.2, and Proposition 3.3
@@
A.4 Derivatives of log-determinants using stochastic trace estimation and variance reduction for SSOR preconditioner
@@
A.4.1 Derivative for variance parameters
@@
A.4.2 Further derivatives for Laplace approximations
@@
A.5 Alternative stochastic trace estimation for Fisher information
@@
$\begin{split}&\operatorname*{tr}\left(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}}\right)\\ =&\operatorname*{tr}\left(\left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{k}Z_{k}^{T}\right.\\ &\left.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{l}Z_{l}^{T}\right)\\ =&\operatorname*{tr}\left(WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}\right)\\ &-2\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}ZW\right)\\ &+\leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}ZW\right)\\ \approx&\operatorname*{tr}\left(WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}\right)\\ &-2\frac{1}{t}\sum_{i=1}^{t}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}ZWP^{-1}z_{i}\\ &+\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{1}{t}\sum_{i=1}^{t}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}WZ_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}ZWP^{-1}z_{i}.\end{split}$ (51)
A.6 Preconditioned conjugate gradient algorithm
@@
1:Matrix $A$, preconditioner matrix $P$, vector $b$
2:$u_{l+1}\approx A^{-1}b$, tridiagonal matrix $\tilde{T}$
@@
4:$\alpha_{0}\leftarrow 1$
5:$\beta_{0}\leftarrow 0$
6:$u_{0}\leftarrow 0$
7:$r_{0}\leftarrow b-Au_{0}$
8:$z_{0}\leftarrow P^{-1}r_{0}$
9:$h_{0}\leftarrow z_{0}$
@@
A.7 Zero fill-in incomplete Cholesky factorization
@@ A.8 Additional results for preconditioner comparison
Figure 9: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when the response variable follows a Bernoulli likelihood with a logit link function. The dashed line represents the result for the Cholesky decomposition.
Figure 10: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when $m_{2}=\frac{m_{1}}{2}$, $m=m_{1}+m_{2}=99^{\prime}999$, $n=10m$, and the response variable follows a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
Figure 11: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when $\sigma^{2}=0.25$, $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, and the response variable follows a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
@@ A.9 Additional results for parameter estimation and prediction in simulated experiments
RMSE $\sigma_{1}^{2}$ $8.97\times 10^{-3}$ $8.96\times 10^{-3}$ $8.83\times 10^{-3}$ $8.83\times 10^{-3}$
Bias $\sigma_{1}^{2}$ $-5.96\times 10^{-4}$ $-6.15\times 10^{-4}$ $-5.57\times 10^{-4}$ $-5.57\times 10^{-4}$
@@
RMSE $\beta_{1}$ $1.17\times 10^{-2}$ $1.17\times 10^{-2}$ $1.17\times 10^{-2}$ $1.17\times 10^{-2}$
Bias $\beta_{1}$ $-2.06\times 10^{-4}$ $-2.06\times 10^{-4}$ $-2.06\times 10^{-4}$ $-2.06\times 10^{-4}$
@@
Bias $\beta_{2}$ $-4.41\times 10^{-4}$ $-4.41\times 10^{-4}$ $-4.41\times 10^{-4}$ $-4.41\times 10^{-4}$
@@
Bias $\beta_{3}$ $-1.14\times 10^{-3}$ $-1.14\times 10^{-3}$ $-1.14\times 10^{-3}$ $-1.14\times 10^{-3}$
@@
RMSE $\sigma_{1}^{2}$ $2.01\times 10^{-2}$ $2.01\times 10^{-2}$ $2.05\times 10^{-2}$
Bias $\sigma_{1}^{2}$ $-1.15\times 10^{-2}$ $-1.15\times 10^{-2}$ $-1.13\times 10^{-2}$
RMSE $\sigma_{2}^{2}$ $2.15\times 10^{-2}$ $2.15\times 10^{-2}$ $2.20\times 10^{-2}$
Bias $\sigma_{2}^{2}$ $-1.36\times 10^{-2}$ $-1.36\times 10^{-2}$ $-1.36\times 10^{-2}$
@@
Bias $\beta_{0}$ $-3.68\times 10^{-4}$ $-3.67\times 10^{-4}$ $-4.76\times 10^{-4}$
@@
Bias $\beta_{2}$ $-6.51\times 10^{-3}$ $-6.53\times 10^{-3}$ $-5.87\times 10^{-3}$
@@
Bias $\beta_{4}$ $-1.24\times 10^{-3}$ $-1.23\times 10^{-3}$ $-1.30\times 10^{-3}$
@@
Bias $\beta_{5}$ $-6.36\times 10^{-3}$ $-6.36\times 10^{-3}$ $-5.61\times 10^{-3}$
@@
$\overline{\text{LS}}$ $-4.01\times 10^{-1}$ $-4.01\times 10^{-1}$
@@
sd($\overline{\text{LS}}$) $1.12\times 10^{-3}$ $1.12\times 10^{-3}$
@@ A.10 Additional results for the real-world applications
Figure 13: Non-zero entries of the matrix $Z^{T}Z$ for the real-world data sets.
@@
cars Cholesky (GPBoost) $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.64\times 10^{-3}$
Krylov (GPBoost) $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.64\times 10^{-3}$
glmmTMB $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.63\times 10^{-3}$
lme4 $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.64\times 10^{-3}$
building_permits Cholesky (GPBoost) $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.01\times 10^{-3}$
Krylov (GPBoost) $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.00\times 10^{-3}$
glmmTMB $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.02\times 10^{-3}$
lme4 $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.01\times 10^{-3}$
@@
upselling Cholesky (GPBoost) $-1.74\times 10^{0}$ $2.32\times 10^{-5}$ $9.42\times 10^{-3}$ $3.86\times 10^{-6}$
Krylov (GPBoost) $-1.74\times 10^{0}$ $2.32\times 10^{-5}$ $9.48\times 10^{-3}$ $3.94\times 10^{-6}$

### v2

@@ A.1 Proofs of results in Section 4
Applying the expansion $(1-x)^{-1}=1+x+x^{2}+O(x^{3})$ with $x=\frac{4}{d}+\epsilon_{m}$ gives (38). Similarly, Theorem A.2, applying the expansion $(1-x)^{-1}=1+x+x^{2}+O(x^{3})$ in the denominator of $\frac{1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}}{1-\frac{2}{\sqrt{d}}-\epsilon_{m}^{\prime}}$, and multiplying by the numerator $1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}$ give (39). ∎
A.2 Proofs of Proposition 3.1, Proposition 3.2, and Proposition 3.3
@@
A.3 Derivatives of log-determinants using stochastic trace estimation and variance reduction for SSOR preconditioner
@@
A.3.1 Derivative for variance parameters
@@
A.3.2 Further derivatives for Laplace approximations
@@
A.4 Alternative stochastic trace estimation for Fisher information
@@
$\begin{split}&\operatorname*{tr}\left(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}}\right)\\ =&\operatorname*{tr}\left(\left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{k}Z_{k}^{T}\right.\\ &\left.\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{l}Z_{l}^{T}\right)\\ =&\operatorname*{tr}\left(WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}\right)\\ &-2\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}ZW\right)\\ &+\penalty 10000\ \penalty 10000\ \operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}ZW\right)\\ \approx&\operatorname*{tr}\left(WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}\right)\\ &-2\frac{1}{t}\sum_{i=1}^{t}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}ZWP^{-1}z_{i}\\ &+\penalty 10000\ \penalty 10000\ \frac{1}{t}\sum_{i=1}^{t}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}WZ_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}ZWP^{-1}z_{i}.\end{split}$ (54)
A.5 Preconditioned conjugate gradient algorithm
@@
1:Matrix AA, preconditioner matrix $P$, vector $b$
[…]
@@
[…]
@@
A.6 Zero fill-in incomplete Cholesky factorization
@@
A.7 Simulating unbalanced random effects designs
The unbalanced designs are obtained as follows. We first generate numbers of repeated occurrences $N_{jk}$ ($\sum_{i=1}^{n}(Z_{k})_{ij}$) of random effect $b_{k,j}$ for level $j$ of component $k$ by simulating from a negative binomial distribution with mean $n/m_{k}-1$ and size parameter $r$ and then adding $1$. The choice ensures that every random effect $b_{k,j}$ occurs at least once, and there is variability in the numbers of repeated occurrences $N_{jk}$. Note that the size, or inverse overdispersion, parameter $r$ control the regularity of the random effects design with smaller $r$’s yielding less regular designs with more variability in the number of occurrences per random effect. Unless stated otherwise, we use a size parameter $r=1$. The resulting total numbers of occurrences $T_{k}=\sum_{j=1}^{m_{k}}N_{jk}$ are typically close to $n$, but not exactly equal. Furthermore, the simulated totals $T_{1}$ and $T_{2}$ for the two random effect components usually differ slightly. To ensure that both components have the same total number of occurrences, we adjust the smaller of the two totals. If, say, the simulated total number of occurrences of the first component is smaller than that of the second component, $T_{1}<T_{2}$, we add $T_{2}-T_{1}$ occurrences to the first component by simulating $T_{2}-T_{1}$ samples from a multinomial distribution with $m_{1}$ factors and probabilities $1/m_{1}$ and adding the resulting count vector to $(N_{1,1},\dots,N_{m_{1},1})$. Once the final values $N_{jk}$ are obtained, we construct the incidence matrices $Z_{k}$ by assigning exactly $N_{jk}$ ones to column $j$ in consecutive rows:
$(Z_{k})_{ij}=1\quad\text{for }i=\Bigl(\sum_{j^{\prime}=1}^{j-1}N_{j^{\prime}k}\Bigr)+1,\dots,\sum_{j^{\prime}=1}^{j}N_{j^{\prime}k},$
and zero otherwise (we use the convention $\sum_{j^{\prime}=1}^{0}N_{j^{\prime}k}=0$). Finally, we randomly permute the rows of $Z_{2}$, which induces a random crossing structure between the two random effect components.
@@ A.8 Additional results for preconditioner comparison
Figure 8: Negative log-likelihood and wall-clock time in seconds for different preconditioners and varying numbers of random vectors $t$ in the SLQ method for a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
Figure 9: Accuracy (in standard deviations; top row), wall-clock time (in seconds; bottom row, left axis), and number of CG iterations (bottom row, right axis) for varying numbers of random vectors $t$ in the SLQ method for calculating log-marginal likelihoods for a Bernoulli likelihood with a logit link.
Figure 10: Negative log-likelihood and wall-clock time in seconds for varying numbers of random vectors $t$ in the SLQ method for a Bernoulli likelihood with a logit link. The dashed line represents the result for the Cholesky decomposition.
Figure 11: Accuracy (in standard deviations; top row), wall-clock time (in seconds; bottom row, left axis), and number of CG iterations (bottom row, right axis) for varying numbers of random vectors $t$ in the SLQ method for calculating log-marginal likelihoods for a Gaussian likelihood with an error variance $\sigma^{2}=0.05$ (= higher signal-to-noise ratio).
Figure 12: Accuracy (in standard deviations; top row), wall-clock time (in seconds; bottom row, left axis), and number of CG iterations (bottom row, right axis) for varying irregularity of the random effects design (left column) and random effects dimension $m$ (right column) for calculating log-marginal likelihoods for a Gaussian likelihood. ‘over_disp’ equals the inverse size parameter $1/r$ in the unbalanced design simulation approach explained in Appendix A.7. A larger ‘over_disp’ value corresponds to a less regular random effects design with more variability in the number of occurrences per random effect. An unbalanced design with $r=1$ is used when varying $m$.
@@ A.9 Additional results for parameter estimation and prediction in simulated experiments
RMSE $\sigma_{1}^{2}$ $8.97\times 10^{-3}$ $96\times 10^{-3} 8.8$$3\times 10^{-3} 8.8$$83\times 10^{-3}$
Bias $\sigma_{1}^{2}$ $5.96\times 10^{-4}$ $6.15\times 10^{-4}$ $57\times 10^{-4} 5.5$$57\times 10^{-4}$
@@
RMSE $\beta_{1}$ $17\times 10^{-2} 1.1$$7\times 10^{-2} 1.1$$7\times 10^{-2} 1.1$$17\times 10^{-2}$
Bias $\beta_{1}$ $2.06\times 10^{-4}$ $2.06\times 10^{-4}$ $2.06\times 10^{-4}$ $2.06\times 10^{-4}$
@@
Bias $\beta_{2}$ $41\times 10^{-4} 4.4$$\times 10^{-4} 4.41×1$$\times 10^{-4} 4.4$$41\times 10^{-4}$
@@
Bias $\beta_{3}$ $14\times 10^{-3} 1.1$$4\times 10^{-3} 1.1$$4\times 10^{-3} 1.1$$14\times 10^{-3}$
@@
RMSE $\sigma_{1}^{2}$ $\times 10^{-2} 2.01×102$$\times 10^{-2}$ $2.05\times 10^{-2}$
Bias $\sigma_{1}^{2}$ $15\times 10^{-2} 1.1$$5\times 10^{-2} 1.1$$13\times 10^{-2}$
RMSE $\sigma_{2}^{2}$ $2.15\times 10^{-2}$ $15\times 10^{-2} 2.2$$20\times 10^{-2}$
Bias $\sigma_{2}^{2}$ $1.36\times 10^{-2}$ $1.36\times 10^{-2}$ $1.36\times 10^{-2}$
@@
Bias $\beta_{0}$ $3.68\times 10^{-4}$ $3.67\times 10^{-4}$ $4.76\times 10^{-4}$
@@
Bias $\beta_{2}$ $6.51\times 10^{-3}$ $6.53\times 10^{-3}$ $5.87\times 10^{-3}$
@@
Bias $\beta_{4}$ $1.24\times 10^{-3}$ $1.23\times 10^{-3}$ $1.30\times 10^{-3}$
@@
Bias $\beta_{5}$ $6.36\times 10^{-3}$ $6.36\times 10^{-3}$ $5.61\times 10^{-3}$
@@
$\overline{\text{LS}}$ $4.01\times 10^{-1}$ $4.01\times 10^{-1}$
@@
sd($\overline{\text{LS}}$) $12\times 10^{-3} 1.1$$12\times 10^{-3}$
@@ A.10 Additional results for the real-world applications
Figure 14: Non-zero entries of the matrix $Z^{T}Z$ for the real-world data sets (except the MovieLens_32m data set for which $Z^{T}Z$ is too large to visualize).
@@
MovieLens_32m Cholesky (GPBoost) crashed
Krylov (GPBoost) 283.8 40613057
glmmTMB crashed
lme4 273152.9 40613057
@@
MovieLens_32m Cholesky (GPBoost) crashed
Krylov (GPBoost) $7.23\times 10^{-1}$ $1.93\times 10^{-1}$ $2.32\times 10^{-1}$
glmmTMB crashed
lme4 $7.23\times 10^{-1}$ $1.93\times 10^{-1}$ $2.32\times 10^{-1}$
@@
cars Cholesky (GPBoost) $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $1.77\times 10^{-1}$ $2.64\times 10^{-3}$
Krylov (GPBoost) $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $1.77\times 10^{-1}$ $2.64\times 10^{-3}$
glmmTMB $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $1.77\times 10^{-1}$ $2.63\times 10^{-3}$
lme4 $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $1.77\times 10^{-1}$ $2.64\times 10^{-3}$
building_permits Cholesky (GPBoost) $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $1.36\times 10^{-5}$ $2.01\times 10^{-3}$
Krylov (GPBoost) $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $1.36\times 10^{-5}$ $2.00\times 10^{-3}$
glmmTMB $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $1.36\times 10^{-5}$ $2.02\times 10^{-3}$
lme4 $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $1.36\times 10^{-5}$ $2.01\times 10^{-3}$
@@
MovieLens_32m Cholesky (GPBoost) crashed
Krylov (GPBoost) $3.39\times 10^{0}$ $7.29\times 10^{-2}$ $4.93\times 10^{-2}$
glmmTMB crashed
lme4 $3.39\times 10^{0}$ $7.29\times 10^{-2}$ $4.93\times 10^{-2}$
@@
upselling Cholesky (GPBoost) $1.74\times 10^{0}$ $2.32\times 10^{-5}$ $9.42\times 10^{-3}$ $3.86\times 10^{-6}$
Krylov (GPBoost) $1.74\times 10^{0}$ $2.32\times 10^{-5}$ $9.48\times 10^{-3}$ $3.94\times 10^{-6}$
glmmTMB crashed
lme4 crashed